OpenStudy (anonymous):

integrate x/sqroot of x^4-1

OpenStudy (anonymous):

nasty integral.

OpenStudy (anonymous):

1/2 cosh^-1(x^2)

OpenStudy (anonymous):

thanks dave,,, but i want to know how to do it

OpenStudy (anonymous):

he made the substitution x = cosh(u)

OpenStudy (anonymous):

step by step

OpenStudy (anonymous):

sorry. he made the substitution x^2 = cosh(u)

OpenStudy (anonymous):

hence 2xdx = sinh(u)du, so the integral reduces to: int(1/2*sinh(u)du/sqrt((cosh(u))^2 - 1)) = int(1/2*sinh(u)du/sinh(u)) = int(1/2 * du) = 1/2 * u + C = 1/2 * arccos(x^2) + C

OpenStudy (anonymous):

\[u=x^2\] The equation now becomes \[1/2\int\limits1/\sqrt(u^2 -1)\] which is the derivative of arcosh(u) Plug the 1/2 back in and switch the u for x^2 and you get \[1/2\cosh^{-1} (x^2)\]